Consider the vectors :

u = (1,0) and v= (0,1)

1) Determine CosA, where A is the angle between u and v.

2) Determine the area of the parallelogram determined by u and v.
1
Expert's answer
2020-04-16T15:15:04-0400

cosA = "\dfrac{u*v}{||u||*||v||}"


u * v = u1v1 + u2v2 = 1*0 + 0*1 = 0+0=0


||u|| = "\sqrt{u_1^2+u_2^2}=\sqrt{1^2+0^2}=\sqrt{1}=1"


||v|| = "\sqrt{v_1^2+v_2^2}=\sqrt{0^2+1^2}=\sqrt{1}=1"


cosA = "\dfrac{0}{1*1}=0"


Area(P) = |u * v|


|u * v| = "\begin{vmatrix}
 u_1 & v_1 \\
 u_2 & v_2
\end{vmatrix}"


|u * v| = "\begin{vmatrix}
 1 & 0 \\
 0 & 1
\end{vmatrix}=u_1*v_2-u_2*v_1=1*1-0*0=1"


Area(P) = 1